OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.

OpenStudy (anonymous):

slope m=y1-y2/x1-x2 = m =(4-3)/(4-1) = 1/3 point slope y-y1=m(x-x1) = y-3=1/3(x-4) = 1/3x-4/3+3 y=1/3x-1/3

OpenStudy (anonymous):

sorry -4/3+3 = 5/3 not -1/3 :/